OpenStudy (axel.caballero):

Derive the equation of the parabola with a focus at (-7, 5) and a directrix of y = -11. Question 2 Derive the equation of the parabola with a focus at (-5, 5) and a directrix of y = -1 Derive the equation of the parabola with a focus at (0, 1) and a directrix of y = -1

OpenStudy (anonymous):

ok so, this is not something that i do often but here goes. do you know the generic equation for a parabola?

OpenStudy (anonymous):

I have to go now so i will leave you with the equation and an example. Generic equation for a parabola:\[(x-h)^{2} = 4p(y-k)\] the coordinates of the vertex are: \[Vertex - (h,k)\] The Coordinates of the vertex are:\[Focus - (h,(k+p))\] The equation for the directix: \[y = k-p\] Therefore your first question - Focus = (-7,5), Directix = -11 From the equation for a directix: \[-11=y=k-p\] From the y coordinate for the focus:\[5 = k-p\] You can then solve these equations simultaneously and get k and p: k = -3, p = 8 you know from the x coodinate of the focus is h = -7 so you can wright the equation for the parabola\[(x-(-7))^{2}=4\times8(y-(-3))\]Simplifies to :\[(x+7)^{2} = 32(y+3)\]